# LD-Scene: LLM-Guided Diffusion for Controllable Generation of Adversarial Safety-Critical Driving Scenarios

### v1

@@ III-B Latent Diffusion Models for Scenario Generation
Diffusion models [36] comprise two Markov chains: a forward (diffusion) process that progressively adds Gaussian noise to the data, leading to pure noise over multiple steps, and a reverse (denoising) process, where a learnable neural network iteratively removes noise to generate samples. In this work, we train a diffusion model to synthesize realistic and controllable adversarial safety-critical driving scenarios through an iterative denoising process.
@@
Unlike conventional diffusion-based trajectory generation methods that operate directly in the trajectory space [40, 41, 18], our approach applies the diffusion model within a latent space. This design choice reduces computational complexity while enhancing the expressiveness of the generative process [22, 9, 16]. Specifically, we integrate the diffusion model with a pretrained graph-based variational autoencoder (VAE) model following Strive [9]. The graph-based VAE consists of an encoder that learns a latent representation of agent interactions and a decoder that autoregressively generates future trajectories using a kinematic bicycle model, ensuring both realism and plausibility in traffic scenario generation.
@@
Architecture. As depicted in the upper section of Fig. 1, our LDM consists of three key components: two encoders with frozen parameters, a learnable denoising network, and an autoregressive trajectory decoder with frozen parameters. The encoders and decoder are directly adopted from a pretrained VAE model, which all these modules employ GNNs. Specifically, the prior scene input encoder $\mathcal{E}_{\theta}$ is identical to the prior network in Strive [9], modeling agent interactions via a fully connected scene graph. Each node encodes contextual features derived from an agent’s past trajectory and local rasterized map. Through message passing, the encoder generates latent representations for all agents, which serve as conditioning inputs for our diffusion model, formulated as $\textit{{c}}=\mathcal{E}_{\theta}(\textit{{x}},\textit{{m}})$. Similarly, the posterior scene input encoder $\mathcal{E}_{\theta}$ corresponds to the posterior network in Strive [9], which operates jointly on past and future information. The resulting latent vectors serve as the latent input $\mathbf{z}$ for our diffusion model, expressed as $\mathbf{z}=\mathcal{E}_{\theta}(\bm{\tau},\textit{{x}},\textit{{m}})$. The diffusion process begins with a clean future latent $\mathbf{z}^{0}\sim q(\mathbf{z}^{0})$ sampled from the data distribution. The forward process produces a sequence of progressively noisier latent $(\mathbf{z}^{0},\mathbf{z}^{1},...,\mathbf{z}^{K})$, where each step $k$ follows a Gaussian noise injection process [36]:
$q(\mathbf{z}^{1:K}\mid\mathbf{z}^{0})$ $:=\prod_{k=1}^{K}q(\mathbf{z}^{k}\mid\mathbf{z}^{k-1})$ (1)
$q(\mathbf{z}^{k}\mid\mathbf{z}^{k-1})$ $:=\mathcal{N}\left(\mathbf{z}^{k};\sqrt{1-\beta_{k}}\mathbf{z}^{k-1},\beta_{k}\mathbf{I}\right)$
@@
The denoising network learns to reverse this noise process. Given a noisy latent input $\mathbf{z}^{K}$, the denoising network $p_{\theta}(\mathbf{z}^{k-1}|\mathbf{z}^{k})$ learns to denoise it that generates $p_{\theta}(\mathbf{z}^{0})$ in a step-by-step manner. Each step in the reverse process is conditioned on c:
$p_{\theta}(\mathbf{z}^{k-1}\mid\mathbf{z}^{k},\textit{{c}}):=\mathcal{N}(\mathbf{z}^{k-1};\mu_{\theta}(\mathbf{z}^{k},k,\textit{{c}}),\Sigma_{k})$ (2)
where $\mu_{\theta}(\cdot)$ is the predicted mean, $\theta$ represents the model parameters, and $\Sigma_{k}$ is a fixed covariance matrix. The model is trained to parameterize the mean of the Gaussian distribution at each step of the denoising process.
As illustrated in Fig. 1, the denoising network adopts a U-Net architecture similar to [39], consisting of a series of one-dimensional temporal convolutional blocks. The conditioning input c is incorporated into the intermediate input latent within each convolutional block.
Finally, the trajectory decoder $\mathcal{D}$, following the architecture in Strive [9], also employs a GNN-based interaction net to ensure realistic vehicle interactions. Specifically, the decoder $\mathcal{D}_{\theta}(\mathbf{z}^{0},\textit{{x}},\textit{{m}})$ operates on the scene graph with both the denoised latent and past information embedding. The decoding process is autoregressive, where the model predicts all agents’ actions $a_{t}$ at timestep $t$, which are then propagated through a kinematic bicycle model $f$ to compute the next state $s_{t+1}$. The updated state is incorporated into the embedding before proceeding to the next time step. This autoregressive formulation ensures the physical plausibility of the generated trajectories while maintaining realistic multi-agent interactions.
Training. Since the VAE is pretrained, we focus on training the LDM. The objective is to train a denoising network that minimizes the variational bound on the negative log-likelihood, which can be simplified to a mean squared error loss between the predicted noise and the actual noise [36]:
@@
III-C Guidance Generation with LLMs
@@
III-C1 User Query and System Prompt
@@
III-C2 Code Generator
CoT reasoning has been shown to be highly effective in problem decomposition and in carrying out more intricate reasoning tasks [51, 52]. Therefore, our code generator adopts CoT prompting to enhance effectiveness and provide greater controllability in adversarial scenario generation. The code generator prompts in our framework consist of two primary modules: the reasoning prompt and the code generation prompt. As shown in Fig. 2(b), the code generation prompt provides a template that ensures the correct implementation of the guidance loss function, while Fig. 2(c) presents the reasoning prompt, which instructs the model to generate the loss function step-by-step.
@@ III-C2 Code Generator
Code generation prompt. Unlike the few-shot learning approach used in CTG++ [41], our method employs a zero-shot learning strategy for code generation. The Code Generation Prompt provides a structured template that directly incorporates examples of some loss functions, including the agent collision loss function and the map collision loss function, as shown in Fig. 2(b). Based on these examples, the LLM learns how to utilize the input trajectory and apply helper classes such as EnvCollLoss and VehCollLoss. Finally, the code generator generates the adversarial collision loss based on the user query. Depending on different user queries, such as cut-in, overtaking, or emergency braking, the corresponding adversarial collision loss function is generated, ensuring controllable adversarial behavior.
III-C3 Code Debugger
As noted in CTG++ [41], code generation can sometimes produce incorrect implementations. To address this limitation, we introduce a code debugger module that reviews and refines the generated code. The debugger operates by running a closed-loop unit test to evaluate the generated guidance loss function. If an error is detected, the corresponding source code and error message are provided to the code debugger, which iteratively refines the implementation until a predefined maximum number of iterations is reached.
@@ III-C3 Code Debugger
III-D Generation of Safety-Critical Scenarios
To enable the controllable generation of safety-critical driving scenarios, we introduce an objective function $\mathcal{J}(\bm{\tau})$, which guides the denoising process. Adversarial driving scenarios are generated by perturbing the predicted mean at each denoising step. Since the denoising process operates in the latent space, each guided iteration step first decodes the latent vector $\mathbf{z}$ into the corresponding trajectory $\bm{\tau}$ using a decoder $\mathcal{D}$, formulated as $\bm{\tau}=\mathcal{D}_{\theta}(\mathbf{z},\textit{{x}},\textit{{m}})$. At each reverse diffusion step $k$, we modify the denoising process by adding the gradient of $\mathcal{J}$ as guidance [39]:
@@ III-D Generation of Safety-Critical Scenarios
To further enhance the robustness of the guidance mechanism, we adopt the reconstruction guidance (clean guidance) strategy [53], which perturbs the clean latent vector $\hat{\mathbf{z}}^{0}$ predicted by the network. This approach mitigates numerical instabilities, ensuring a more stable and controllable denoising process.
In detail, the objective function $\mathcal{J}(\bm{\tau})$ comprises three components: (i) $\mathcal{J}_{bv\_real}$, which ensures that non-adversarial vehicles behave realistically by preventing collisions with each other and avoiding off-road deviations; (ii) $\mathcal{J}_{adv\_real}$, which maintains the plausibility of the adversarial vehicle’s behavior by preventing it from colliding with non-adversarial vehicles or leaving the roadway; and (iii) $\mathcal{J}_{adv}$, which is generated by the LLM mentioned in the previous section and formulated to control the adversarial vehicle’s behavior to induce a collision with the ego vehicle based on a user query. By integrating these objectives, the proposed framework effectively balances realism while enabling the controllable generation of adversarial driving scenarios.
@@
[…]
@@
[…]
@@ IV Experimental Results
This section outlines the evaluations of our proposed LD-Scene model. We first describe the dataset, evaluation metrics, and experimental setup. We then compare the performance of our LD-Scene against baseline models and provide a quantitative analysis of the results. Additionally, we conduct an ablation study to assess the contributions of key components in our method. Finally, we perform two controllability studies, examining both the controllable adversarial level and controllable adversarial behavior.
IV-A Dataset
@@
We conduct our experiments on the nuScenes [54] dataset, which comprises 1,000 driving scenes, each lasting 20 seconds and recorded at 2 Hz. The dataset captures 5.5 hours of urban driving data collected from two cities, Boston and Singapore, covering diverse traffic conditions and complex interactions between road agents. We train our models using the training split and evaluate them on the validation split of nuScenes. Following the standard guidelines of the nuScenes prediction challenge, we use 2 seconds (4 steps) of past motion data to predict the future 6 seconds (12 steps) of trajectories.
IV-B Evaluation Metrics
@@
We focus on realistic safety-critical scenarios and propose a set of metrics to evaluate the quality of the generated scenarios, mainly in three aspects: adversariality, realism, and diversity.
@@
  • Adversariality: This aspect measures the effectiveness of the generated scenarios in creating safety-critical situations for the ego vehicle. Adv-Ego Collision Rate quantifies the percentage of scenarios where the adversarial vehicle collides with the ego vehicle, with higher values indicating stronger adversarial effectiveness. Adv Acceleration (Adv Acc) measures the acceleration magnitude of the adversarial vehicle, where higher values represent more aggressive adversarial behavior.
@@
  • Realism: To ensure the generated scenarios align with real-world traffic behavior, we assess both offroad rates and collision rates. Adv Offroad Rate and Other Offroad Rate measure the frequency at which adversarial and non-adversarial agents drive offroad, where lower values indicate more plausible behavior. Collision rates, including Adv-Other Coll Rate, Other-Ego Coll Rate, and Other-Other Coll Rate, quantify the frequency of crashes among different vehicle groups, with lower values ensuring realistic interactions. Additionally, Other Acceleration (Other Acc) measures the acceleration magnitude of non-adversarial agents, where lower values suggest smoother and more realistic driving behavior.
  • Diversity: To evaluate the diversity of the generated scenarios, we use the Final Displacement Diversity (FDD) metric, following [31, 16], to measure the maximum difference in final displacement among variations. A higher FDD value indicates greater scenario diversity and broader coverage of driving scenarios.

### v2

@@ 3.2 Latent Diffusion Models for Scenario Generation
Diffusion models (Ho et al., 2020) comprise two Markov chains: a forward (diffusion) process that progressively adds Gaussian noise to the data, leading to pure noise over multiple steps, and a reverse (denoising) process, where a learnable neural network iteratively removes noise to generate samples. In this work, we train a diffusion model to synthesize realistic and controllable adversarial safety-critical driving scenarios through an iterative denoising process.
@@
Unlike conventional diffusion-based trajectory generation methods that operate directly in the trajectory space (Zhong et al., 2023b, a; Chang et al., 2024), our approach applies the diffusion model within a latent space. This design choice reduces computational complexity while enhancing the expressiveness of the generative process (Rombach et al., 2022; Rempe et al., 2022; Xie et al., 2024). Specifically, we integrate the diffusion model with a pretrained graph-based variational autoencoder (VAE) model following Strive (Rempe et al., 2022). The graph-based VAE consists of an encoder that learns a latent representation of agent interactions and a decoder that autoregressively generates future trajectories using a kinematic bicycle model, ensuring both realism and plausibility in traffic scenario generation.
@@
Architecture. As depicted in the upper section of Fig. 1, our LDM consists of three key components: two encoders with frozen parameters, a learnable denoising network, and an autoregressive trajectory decoder with frozen parameters. The encoders and decoder are directly adopted from a pretrained VAE model, which all these modules employ GNNs. Specifically, the prior scene input encoder $\mathcal{E}_{\theta}$ is identical to the prior network in Strive (Rempe et al., 2022), modeling agent interactions via a fully connected scene graph. Each node encodes contextual features derived from an agent’s past trajectory and local rasterized map. Through message passing, the encoder generates latent representations for all agents, which serve as conditioning inputs for our diffusion model, formulated as $\textit{{c}}=\mathcal{E}_{\theta}(\textit{{x}},\textit{{m}})$. Similarly, the posterior scene input encoder $\mathcal{E}_{\theta}$ corresponds to the posterior network in Strive (Rempe et al., 2022), which operates jointly on past and future information. The resulting latent vectors serve as the latent input $\mathbf{z}$ for our diffusion model, expressed as $\mathbf{z}=\mathcal{E}_{\theta}(\boldsymbol{\tau},\textit{{x}},\textit{{m}})$. The diffusion process begins with a clean future latent $\mathbf{z}^{0}\sim q(\mathbf{z}^{0})$ sampled from the data distribution. The forward process produces a sequence of progressively noisier latent $(\mathbf{z}^{0},\mathbf{z}^{1},...,\mathbf{z}^{K})$, where each step $k$ follows a Gaussian noise injection process (Ho et al., 2020):
$q(\mathbf{z}^{1:K}\mid\mathbf{z}^{0})$ $=\prod_{k=1}^{K}q(\mathbf{z}^{k}\mid\mathbf{z}^{k-1})$ (1)
$q(\mathbf{z}^{k}\mid\mathbf{z}^{k-1})$ $=\mathcal{N}\left(\mathbf{z}^{k};\sqrt{1-\beta_{k}}\mathbf{z}^{k-1},\beta_{k}\mathbf{I}\right)$
@@
To accelerate the inference process, we employ the Denoising Diffusion Implicit Models (DDIM) sampling strategy (Song et al., 2020), which facilitates a non-Markovian formulation of the reverse diffusion process. Unlike conventional sampling schemes, DDIM enables efficient generation by skipping intermediate denoising steps, thus significantly reducing the number of sampling iterations without necessitating model retraining. The reverse process at step $k$ is defined as:
$\mathbf{z}^{k-1}=\sqrt{\alpha_{k-1}}\cdot\tilde{\mathbf{z}}^{0}+\sqrt{1-\alpha_{k-1}}\cdot\epsilon_{\theta}(\mathbf{z}^{k},k,\mathbf{c})$ (2)
where $\epsilon_{\theta}(\mathbf{z}^{k},k,\textit{{c}})$ denotes the denoising network that predicts the noise conditioned on the step $k$ and conditioning latent c. The coefficient $\alpha_{k}=\prod_{i=1}^{k}(1-\beta_{i})$ represents the cumulative product of the noise schedule. The estimate of the clean latent representation $\tilde{\mathbf{z}}^{0}$ at step $k$ is given by:
$\tilde{\mathbf{z}}^{0}=\left(\frac{\mathbf{z}^{k}-\sqrt{1-\alpha_{k}}\cdot\epsilon_{\theta}(\mathbf{z}^{k},k,\mathbf{c})}{\sqrt{\alpha_{k}}}\right)$ (3)
By iteratively applying this reverse denoising operation starting from the noisy latent $\mathbf{z}^{K}$, we obtain the final denoised sample $\hat{\mathbf{z}}^{0}$.
As illustrated in Fig. 1, the denoising network adopts a U-Net architecture similar to Janner et al. (2022), composed of one-dimensional temporal convolutional blocks with stacked residual blocks and linear attention. The conditioning input c is incorporated into the intermediate input latent within each convolutional block.
Finally, the trajectory decoder $\mathcal{D}$, following the architecture in Strive (Rempe et al., 2022), also employs a GNN-based interaction net to ensure realistic vehicle interactions. Specifically, the decoder $\mathcal{D}_{\theta}(\hat{\mathbf{z}}^{0},\textit{{x}},\textit{{m}})$ operates on the scene graph with both the denoised latent and past information embedding. The decoding process is autoregressive, where the model predicts all agents’ actions $a_{t}$ at timestep $t$, which are then propagated through a kinematic bicycle model $f$ to compute the next state $s_{t+1}$. The updated state is incorporated into the embedding before proceeding to the next time step. This autoregressive formulation ensures the physical plausibility of the generated trajectories while maintaining realistic multi-agent interactions.
Training. Since the VAE is pretrained, we focus on training the LDM. The objective is to train a denoising network that minimizes the variational bound on the negative log-likelihood, which can be simplified to a mean squared error loss between the predicted noise and the actual noise:
@@
3.3 Guidance Generation with LLMs
@@
3.3.1 User Query and System Prompt
@@
3.3.2 Code Generator
CoT reasoning has been shown to be highly effective in problem decomposition and in carrying out more intricate reasoning tasks (Wei et al., 2022; Kojima et al., 2022). Therefore, our code generator adopts CoT prompting to enhance effectiveness and provide greater controllability in adversarial scenario generation. The code generator prompts in our framework consist of two primary modules: the reasoning prompt and the code generation prompt. As shown in Fig. 2(b), the code generation prompt provides a template that ensures the correct implementation of the guidance loss function, while Fig. 2(c) presents the reasoning prompt, which instructs the model to generate the loss function step-by-step.
@@ 3.3.2 Code Generator
Code generation prompt. Unlike the few-shot learning approach used in CTG++ (Zhong et al., 2023a), our method employs a zero-shot learning strategy for code generation. The Code Generation Prompt provides a structured template that directly incorporates examples of some loss functions, including the agent collision loss function and the map collision loss function, as shown in Fig. 2(b). Based on these examples, the LLM learns how to utilize the input trajectory and apply helper classes such as EnvCollLoss and VehCollLoss. Finally, the code generator generates the adversarial collision loss based on the user query. Depending on different user queries, such as cut-in, overtaking, or emergency braking, the corresponding adversarial collision loss function is generated, ensuring controllable adversarial behavior.
3.3.3 Code Debugger
As noted in CTG++ (Zhong et al., 2023a), code generation can sometimes produce incorrect implementations. To address this limitation, we introduce a code debugger module that reviews and refines the generated code. The debugger operates by running a closed-loop unit test to evaluate the generated guidance loss function. If an error is detected, the corresponding source code and error message are provided to the code debugger, which iteratively refines the implementation until a predefined maximum number of iterations is reached.
@@ 3.3.3 Code Debugger
3.4 Generation of Safety-Critical Scenarios
To enable the controllable generation of safety-critical driving scenarios, we introduce an objective function $\mathcal{J}(\boldsymbol{\tau})$, which guides the denoising process. Adversarial driving scenarios are generated by perturbing the predicted mean at each denoising step. Since the denoising process operates in the latent space, each guided iteration step first decodes the latent vector $\mathbf{z}$ into the corresponding trajectory $\boldsymbol{\tau}$ using a decoder $\mathcal{D}$, formulated as $\boldsymbol{\tau}=\mathcal{D}_{\theta}(\mathbf{z},\textit{{x}},\textit{{m}})$. At each reverse diffusion step $k$, we modify the denoising process by adding the gradient of $\mathcal{J}$ as guidance (Janner et al., 2022):
@@ 3.4 Generation of Safety-Critical Scenarios
To further enhance the robustness of the guidance mechanism, we adopt the reconstruction guidance (clean guidance) strategy (Rempe et al., 2023), which perturbs the clean latent vector $\hat{\mathbf{z}}^{0}$ predicted by the network. This approach mitigates numerical instabilities, ensuring a more stable and controllable denoising process.
In detail, the objective function $\mathcal{J}(\boldsymbol{\tau})$ comprises three components: (i) $\mathcal{J}_{bv\_real}$, which ensures that non-adversarial vehicles behave realistically by preventing collisions with each other and avoiding off-road deviations; (ii) $\mathcal{J}_{adv\_real}$, which maintains the plausibility of the adversarial vehicle’s behavior by preventing it from colliding with non-adversarial vehicles or leaving the roadway; and (iii) $\mathcal{J}_{adv}$, which is generated by the LLM mentioned in the previous section and formulated to control the adversarial vehicle’s behavior to induce a collision with the ego vehicle based on a user query. By integrating these objectives, the proposed framework effectively balances realism while enabling the controllable generation of adversarial driving scenarios.
@@
$\text{veh\_coll\_pens}_{ij}(t)=\begin{cases}1-\frac{d_{ij}(t)}{p_{ij}},&\text{if }d_{ij}(t)\leq p_{ij}\\ 0,&\text{otherwise}\end{cases}$ (6)
@@
$\text{env\_coll\_pens}_{i}(t)=\begin{cases}1-\frac{d_{i}(t)}{p_{i}},&\text{if }d_{i}(t)\leq p_{i}\\ 0,&\text{otherwise}\end{cases}$ (7)
@@ 4 Experimental Results
This section outlines the evaluations of our proposed LD-Scene model. We first describe the dataset, evaluation metrics, and experimental setup. We then compare the performance of our LD-Scene against baseline models and provide a quantitative analysis of the results. Additionally, we perform two ablation studies: one to demonstrate the effectiveness of the guidance components and another to validate the effectiveness of the debugger module. Finally, we perform two controllability studies, examining both the controllable adversarial level and controllable adversarial behavior.
4.1 Dataset
@@
We conduct our experiments on the nuScenes (Caesar et al., 2020) dataset, which comprises 1,000 driving scenes, each lasting 20 seconds and recorded at 2 Hz. The dataset captures 5.5 hours of urban driving data collected from two cities, Boston and Singapore, covering diverse traffic conditions and complex interactions between road agents. We train our models using the training split and evaluate them on the validation split of nuScenes. Following the standard guidelines of the nuScenes prediction challenge, we use 2 seconds (4 steps) of past motion data to predict the future 6 seconds (12 steps) of trajectories.
4.2 Evaluation Metrics
@@
We focus on realistic safety-critical scenarios and propose a set of metrics to evaluate the quality of the generated scenarios, mainly in three aspects: adversariality, Behavior Plausibility, and efficiency.
@@
  • Adversariality: This aspect measures the effectiveness of the generated scenarios in simulating safety-critical situations for the ego vehicle. Adv-Ego Collision Rate quantifies the percentage of scenarios where the adversarial vehicle collides with the ego vehicle, with higher values indicating stronger adversarial effectiveness. Adv Acceleration (Adv Acc) measures the acceleration magnitude of the adversarial vehicle, where higher values represent more aggressive adversarial behavior.
@@
  • Behavior Plausibility: To ensure the generated scenarios align with real-world traffic behavior, we assess both offroad rates and collision rates. Adv Offroad Rate and Other Offroad Rate measure the frequency at which adversarial and non-adversarial agents drive offroad, where lower values indicate more plausible behavior. Collision rates, including Adv-Other Coll Rate, Other-Ego Coll Rate, and Other-Other Coll Rate, quantify the frequency of crashes among different vehicle groups, with lower values ensuring realistic interactions.
  • Efficiency: We use the closed-loop simulation time (Sim Time) as the metric, where a shorter simulation time indicates higher efficiency.
4.3 Experimental Setup
Baseline. We evaluate our proposed method against existing baseline approaches for generating adversarial driving scenarios. Specifically, we compare with: our re-implementation of AdvSim (Wang et al., 2021), which optimizes the acceleration of a predefined adversarial vehicle to induce a collision, with initial states generated by SimNet (Bergamini et al., 2021); and Strive (Rempe et al., 2022), for which we utilize its open-source implementation. Strive employs a learned traffic model and performs adversarial optimization in the latent space. We further compare with two diffusion-based baselines: DiffScene (Xu et al., 2023), which incorporates a human-designed safety guidance function to guide the generation toward safety-critical scenarios; and Safe-Sim (Chang et al., 2024), which enhances the guidance function by explicitly incorporating time-to-collision (TTC) constraints to generate more targeted adversarial scenarios.
Model Adversariality Behavior Plausibility Efficiency
Adv-Ego Coll (%) $\uparrow$ Adv Acc (m/s2) $\uparrow$ Adv Offroad (%) $\downarrow$ Other Offroad (%) $\downarrow$ Adv-Other Coll (%) $\downarrow$ Other-Ego Coll (%) $\downarrow$ Other-Other Coll (%) $\downarrow$ Sim Time (s) $\downarrow$
AdvSim 24.72 0.90 15.60 14.85 0.56 0.91 0.11 338.35
Strive 22.69 0.88 18.94 16.64 0.90 1.08 0.05 609.72
DiffScene 15.06 0.98 19.71 19.65 8.03 2.60 1.67 199.01
Safe-Sim 27.81 1.09 21.79 18.12 7.52 3.21 0.66 193.59
LD-Scene 40.75 1.36 12.52 17.95 4.93 2.17 0.66 229.40
Table 1: Overall performance comparison of baseline models on the nuScenes dataset. We compare our approach against AdvSim (Wang et al., 2021), Strive (Rempe et al., 2022), DiffScene (Xu et al., 2023), and Safe-Sim (Chang et al., 2024) for closed-loop safety-critical traffic simulation with a rule-based planner.
Implementation Details. Our LD-Scene is implemented using the PyTorch framework and trained on four GeForce RTX 4090 GPUs for six hours. The diffusion model is trained for 200 epochs using the Adam optimizer with a learning rate of $5\times 10^{-4}$. The number of diffusion steps is set to 20. During the inference stage, we generate multiple candidate future trajectories for each non-ego agent in a given scene, with the number of test samples set to 10. The final trajectory is selected as the one that minimizes the guidance loss and simultaneously satisfies the physical feasibility constraints (Peng et al., 2025). Additionally, both the code generator and debugger used in generating guidance utilize the GPT-4o model.
To ensure a fair evaluation of the controllability of the generated safety-critical scenarios, we adopt a standardized strategy for selecting the adversarial vehicle for each model. Specifically, we assign the adversarial vehicle as the one closest to the ego vehicle in the initial state of the scenario and ensure that it satisfies the required feasibility conditions. This selection strategy follows the approach used in Strive (Rempe et al., 2022); however, unlike Strive, the adversarial vehicle remains fixed throughout the scenario and does not change dynamically. Moreover, all experiments are conducted under a closed-loop simulation setup, where the ego vehicle is controlled by a rule-based lane-graph planner (Montemerlo et al., 2008), ensuring consistent and reactive behavior throughout the scenario.
Figure 3: Visualizations of overall performance among scenario generation models. (a)–(d): Distributions of longitudinal/lateral speed and acceleration. (e): Radar chart comparing adversariality, behavior plausibility, and generation efficiency. LD-Scene achieves the best overall balance.
4.4 Overall Performance
The quantitative performance of our approach and the baselines is presented in Table 1. Compared to the baselines, the proposed LD-Scene demonstrates significant advantages in adversarial metrics while maintaining realistic and plausible agent behaviors. Specifically, our LD-Scene achieves an Adv-Ego Coll (%) of 40.75%, which is substantially higher than the baseline models. This highlights the superior capability of our model in controllably generating safety-critical driving scenarios that involve adversarial vehicles colliding with the ego vehicle. In terms of realism, the LD-Scene shows a lower Adv Offroad (%) of 12.52% compared to AdvSim (15.60%), Strive (18.94%), DiffScene (19.71%), and Safe-Sim (21.79%), indicating our model can better generate adversarial scenarios that are realistic and feasible within the road constraints. The LD-Scene also maintains an overall good performance across other metrics in realism. For instance, the Other Offroad (%) is 17.95%, which is comparable to the baseline models. Additionally, the collision rates involving other vehicles (Adv-Other Coll, Other-Ego Coll, and Other-Other Coll) are all within reasonable ranges, further confirming the realistic nature of the generated scenarios. Furthermore, our diffusion-based framework improves generation efficiency, with LD-Scene achieving an average inference time of 229.40 seconds, significantly faster than test-time optimization methods.
Moreover, Fig. 3(a)-(d) illustrate the driving behavior distributions of non-adversarial agents. Our LD-Scene exhibits more concentrated and narrower distributions, indicating smoother and more stable driving behavior compared to the baseline models. In particular, the narrower lateral speed and acceleration distributions show that LD-Scene avoids unreasonable sharp turns, while the compact longitudinal distributions reflect stable speed regulation. Fig. 3(e) presents a radar chart summarizing each model’s performance in adversariality, behavior plausibility, and efficiency. Here, LD-Scene achieves the highest adversariality score, demonstrating its superior ability to induce safety-critical events, and also records strong behavior plausibility and efficiency values. This balanced performance confirms that LD-Scene achieves the strongest adversarial effectiveness while maintaining high levels of realism and generation efficiency, thus demonstrating its superiority across all evaluation metrics.
Overall, these results demonstrate the capability of our framework to generate realistic safety-critical driving scenarios, which can serve as valuable stress tests for autonomous vehicle systems.